OpenStudy (anonymous):

5. use mental math to determine the solution to x - 3 = 10 ( 1point ) A. x=7 b. x=13 c.x=30 d.x=0

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Add 3 to both sides: x -3 + 3 = 10 + 3 x + 0 = 10 + 3 x = 10 + 3 x = 13

OpenStudy (anonymous):

correct